OpenStudy (anonymous):

Look at the formula for the volume of a cylinder below. V = pi r2h Which of the following equations can be used to find the height (h)?

OpenStudy (anonymous):

\[V=\pi r2h\]

OpenStudy (sriram):

h=V/pi r^2

OpenStudy (anonymous):

h=v/ pi r^2

OpenStudy (anonymous):

\(V=\pi r^2h \implies h=\frac{V}{\pi r^2}\).

OpenStudy (anonymous):